OpenStudy (anonymous):

help will fan and medal A coin bank contains nickels, dimes and quarters totaling $5.45. If there are twice as many quarters as dimes and 11 more nickels than quarters, how many of each coin are in the bank? a. 15 nickels, 2 dimes, and 4 quarters b. 25 nickels, 7 dimes, and 14 quarters c. 17 nickels, 3 dimes, and 6 quarters d. 31 nickels, 5 dimes, and 10 quarters

OpenStudy (anonymous):

Its Option B 25nickels 7dimes and 14quarters

OpenStudy (anonymous):

Take no. of nickels = n dimes = d and quarters = q we have q=2d and n=11+2d so And now d=10 cents q=25cents and n=5cents so 10d+5n+25q=$5.45=545cents or 10d+5(11+2d)+25x2d=545 or 70d+55=545 or 70d=490 or d=7 so q=2x7=14 and n=11+q=25 Yeah Done